OpenStudy (anonymous):

Which of the following equations has only one solution? x 2 + 4x + 4 = 0 x 2 + x = 0 x 2 - 1 = 0

OpenStudy (anonymous):

the one that is a perfect square

jhonyy9 (jhonyy9):

what do you know from dixcriminant property ?

OpenStudy (anonymous):

the second one is a perfect square

jhonyy9 (jhonyy9):

can you answer my question please ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

no actually the second one is not a perfect square \[x^2+4x+4=(x+2)^2\] \[x^2+x=x(x+1)\] \[x^2-1=(x+1)(x-1)\]

jhonyy9 (jhonyy9):

how do you know when the discriminant of one equation is equal zero what sign this for sulutions of this equation ?